A chemist weighed out 69.7g of gallium. Calculate the number of moles of gallium she weighed out


1
Expert's answer
2021-02-24T03:37:37-0500

Standard atomic weight (Ar) of Gallium = 69.7 g/mol


n(Ga)=m/Ar=69.7/69.7=1 mol;


Answer: The chemist weighed out 1 mol of gallium